The equation (sin x)^6 +(cos x)^6=a has a real solution for which of the below mentioned cases:-
A.0.5<=a<=1
B.0.25<=a<=0.5
C.-1<=a<=1
D.0<=a<=0.5

Answer by venugopalramana(3286)

The equation (sin x)^6 +(cos x)^6=a
A={(SINX)^2+(COSX)^2}^3-3*((SINX)^2)*((COSX)^2)*((SINX)^2+(COSX)^2)
=1-3*((SINX)^2)*((COSX)^2)=1-(3/4)*{(SIN(2X))^2
=1-0=1...MAXIMUM......AND.........
=1-1*3/4=0.25...MINIMUM
ANSWER IS A AND B